FriskIsNice:

And this? I’m so so sorry I need help on all 6 questions, I just don’t understand 😢

FriskIsNice:

pooja195:

They want us to solve for D so what operation would you perform to get D alone?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 They want us to solve for D so what operation would you perform to get D alone? \(\color{#0cbb34}{\text{End of Quote}}\) I think subtract -5.004 - 2.826?

pooja195:

Ok lets try to rephrase: how can we can ride of the 5.004 from the left side?

pooja195:

If it is a negative number how can we cancel it out

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 If it is a negative number how can we cancel it out \(\color{#0cbb34}{\text{End of Quote}}\) Add a positive?

pooja195:

Yes so in that case how would you get d by itself?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Yes so in that case how would you get d by itself? \(\color{#0cbb34}{\text{End of Quote}}\) Add -5.004 and 2.826 to find out what d is?

pooja195:

Treat the problem like so: D-5.004=2.826

pooja195:

Disregard the plus sign

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Disregard the plus sign \(\color{#0cbb34}{\text{End of Quote}}\) Umm... add D into 2.826?

pooja195:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Treat the problem like so: D-5.004=2.826 \(\color{#0cbb34}{\text{End of Quote}}\) If we want to cancel out 5.004 (not -5.004) what would we do?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 \(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Treat the problem like so: D-5.004=2.826 \(\color{#0cbb34}{\text{End of Quote}}\) If we want to cancel out 5.004 (not -5.004) what would we do? \(\color{#0cbb34}{\text{End of Quote}}\) Add a positive so make 5.004 + d? Idk I didn’t get a good grade at this

pooja195:

We want to get d alone so 2.826+5.004=?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 We want to get d alone so 2.826+5.004=? \(\color{#0cbb34}{\text{End of Quote}}\) 7.83

pooja195:

Yes so in the future look at it like this: x-10=20 We would add 10 to both sides to get x alone X=30 Just an example

FriskIsNice:

Okay and I got this far and I’m wondering if this is right.